OpenStudy (anonymous):

need help k^3-k + (k+2)!/(k-1)! = (k+1)^3-k+1

OpenStudy (anonymous):

\[k^3-k+\frac{(k+2)!}{(k-1)!}\]??

OpenStudy (anonymous):

or is is all in the numerator?

OpenStudy (anonymous):

is a problem of mathematical induction

OpenStudy (anonymous):

yeah i figured that i was just trying to ascertain if i wrote it correctly

OpenStudy (anonymous):

yes u wrote it correctly

OpenStudy (anonymous):

ok then do you know how to reduce \[\frac{(k+2)!}{(k-1)!}\]

OpenStudy (anonymous):

thats my problem

OpenStudy (anonymous):

i forgot how this so looks familiar thou

OpenStudy (anonymous):

if you wrote out the first few terms it will be completely obvious

OpenStudy (anonymous):

ok ill write it out

OpenStudy (anonymous):

\[(k+2)!=(k+2)(k+1)k(k-1)(k-2)...\]

OpenStudy (anonymous):

is it clear that everything cancels except the first three terms?

OpenStudy (anonymous):

yes we use the bottom as the end

OpenStudy (anonymous):

so only the end cancels out

OpenStudy (anonymous):

so then i can make them equal by rearranging them right